# Cooperation as Black Box: Conceptual Fluctuation and Diagnostic Tools for Misalignment in MAS

### v1

@@ 3 The Misalignment Mosaic
In MAS design, misalignment can occur not only from agent behavior itself, but from gaps between layers of design, implementation, and interpretation. Assumptions made early in the modeling process (often informal, idealized, or morally charged) get embedded in the architecture, translated into behavior, and eventually perceived by stakeholders as intentional system properties. This layered fluctuation creates a critical question: where exactly did the misalignment originate? And who is accountable for its consequences?
@@
Motivated by addressing three interconnected drivers of interpretive misalignment in MAS, the Misalignment Mosaic offers a framework for uncovering the above-mentioned layers. It details four key components contributing to meaning-level misalignment in MAS, depicted in Figure 1, and detailed in the next subsections. It provides a structured diagnostic framework to trace how misalignment emerges from language, framing, and epistemic assumptions, beyond technical mechanisms alone. Rather than resolve ambiguity through formal redefinition, the Mosaic enables audits of meaning, illuminating where alignment falters at the conceptual level.
@@
This framework builds on the concept of alignment magnification, which raises a critical question: How can AI systems be designed to pursue goals that remain consistent with their designers’ intentions? Dung [11] warns that as AI systems become more capable, misalignment risks increase, making these systems not only more dangerous but potentially harder to align. Therefore, alignment should be considered foundational in developing MAS: “No agent exists in a vacuum, and the evaluation of ethical behavior is a complex social and temporal phenomenon” [36].
@@
In fact, the Morality-as-Cooperation Theory serves as an urgent reminder that treating cooperation as a black box is a key source of misalignment in MAS; this conflation is not just a semantic detail, but a structural risk factor in how agent behavior is interpreted, evaluated, and morally framed. Addressing alignment must, therefore, proceed in parallel with refining our understanding of coordination and cooperation (along with their variants). The Misalignment Mosaic integrates seemingly distinct issues to show their role in challenging alignment in MAS. When system designers fail to address ambiguities, agents may operate under misaligned assumptions, which can lead to goal failures, inefficiencies, or emergent behaviors that contradict human intent. Clarifying these concepts is essential for designing MAS systems that operate reliably within both technical and societal constraints.
To scaffold the Mosaic’s utility in MAS design and evaluation workflows, we will develop tools in future work that extend its conceptual framework into practical reflection instruments. These tools will be inspired by the approach of Winikoff et al. [37], who introduced a Scoresheet focused on explainability. Their tool, applicable to MAS and broader AI technologies, incorporates stakeholder requirements and complements the IEEE P7001 Scoresheet [38], which focuses on transparency.
@@ 3.1 Terminological Inconsistency
The coordination/cooperation ambiguity (see section 2) leads to terminological inconsistencies, where these concepts are often defined interchangeably or imprecisely. This fragmentation hinders theoretical and empirical progress, making it difficult to compare findings. The fact that cooperation is still inside a black box shows no surprise that design patterns for MAS are not very popular.
A review of the design patterns literature [39] reveals a notable absence of standardized templates, significantly impacting MAS practical application. This lack of uniformity results in inconsistent implementations, misinterpretations, and reduced transferability across systems. Furthermore, according to Juziuk and colleagues [39], the relationships between patterns are poorly articulated, leading to a fragmented understanding of the pattern space: experts struggle to navigate and apply patterns effectively, limiting their practical utility. While MAS design patterns have been applied across various domains, existing classifications are largely confined to isolated catalogs, which prevents broader integration and adaptation [39]. A key example is the lack of standardized documentation and taxonomies for coordination and cooperation design patterns (along with variants). Without well-established standards, researchers apply similar terms to structurally distinct agent designs. While this issue also has interpretive consequences, we separate it from broader Interpretive Ambiguity by anchoring it in terminological and design-level fragmentation. As MAS applications extend to large-scale deployment across critical industries such as healthcare, transportation systems, and food and agriculture [40], the demand for systems that are not only efficient but also transparent and ethically grounded intensifies.
@@ 3.2 Interpretive Ambiguity
Illustrated by the Rabbit-Duck illusion, this component highlights ambiguities and interpretational limits in MAS research. Just as perception shifts in the illusion, MAS frameworks can be framed differently depending on theoretical or empirical perspectives, leading to inconsistent definitions of coordination and cooperation. In complex or uncertain environments, incomplete knowledge and partial observability further blur the distinction between coordinating tasks and cooperating for collective benefit, amplifying the Misalignment Mosaic. While the Rabbit–Duck illusion broadly represents perceptual ambiguity, we use it here to illustrate interpretive shift between cooperation and coordination (and related concepts); a shift that has conceptual and moral consequences in MAS design.
Key issues can arise when AI system goals are inferred solely from observations, potentially leading to misinterpretations of both the system’s behavior and its intended objectives. Thus, ambiguity in defining cooperation frequently emerges in practice: consider agricultural robots tasked with harvesting crops. At first glance, their interactions may appear cooperative since they collectively contribute to the same task. However, closer analysis may reveal only coordination, where robots follow predefined movement patterns to avoid overlap rather than actively adapting to each other’s presence or modifying strategies based on shared goals. Or, in another example, if various self-driving cars approach a four-way stop, should they coordinate by following predefined traffic rules or activate a cooperative mechanism that helps them dynamically yield? This ambiguity extends beyond system behavior to expert interpretation. Just as in the Rabbit-Duck illusion, it is possible that two researchers analyzing the same MAS testbed reach conflicting conclusions: one classifying it as a coordinative system, while the other sees cooperative behavior. This raises key epistemic and technical questions:
@@
As we try to address the questions above, we observe that multiple factors systematically influence whether MAS behavior is interpreted as coordinative or as indicative of cooperation. These factors span at least three levels (note that [22] provides insights into those):
@@
If Coordination Exists, Is Cooperation Needed?
@@
Klassen et al. [41] demonstrate that partial knowledge can significantly impact agent beliefs and decision-making, sometimes leading to undesirable or even catastrophic outcomes. This raises a critical question for MAS: when agents face epistemic uncertainty, is coordination alone sufficient, or is cooperation necessary to mitigate these risks? In real-world deployments, agents rarely possess complete contextual knowledge due to limitations in data acquisition, processing, and environmental unpredictability. While coordination alone may suffice in simpler, well-defined settings, real-world challenges require strategies that extend beyond simple coordination. Agents must navigate uncertainty, evolving conditions, and imperfect information, making cooperation a potentially necessary mechanism for adaptation and robustness. Given interpretive ambiguity, agents cannot rely solely on their knowledge and beliefs, as these limitations can adversely affect their decision-making capabilities (epistemic side effects [41]). This raises yet another question: if agents had perfect knowledge, would coordination alone be sufficient for optimal decision-making?
@@
While theoretically possible, agents will always operate under knowledge constraints in practice. Given this reality, could a cooperative model serve as a secondary mechanism, allowing agents to compensate for knowledge gaps and make more reliable decisions? This aligns with Curry et al. [42], who demonstrate that cooperative frameworks, grounded in shared goals and moral principles, offer effective approaches for addressing uncertainty and achieving collective success. Incorporating cooperative mechanisms into MAS may thus provide a more effective means of addressing epistemic limitations, ensuring more resilient decision-making in uncertain environments.
@@
Focusing on distributed agent systems, Huhns and Stephens [22] treat agents collectively (as a society of agents) to address scalability challenges in environments with large agent populations. Their work presents a systematic approach to analyzing, describing, and designing multi-agent environments, taking into account both agent topology and content, which evolve dynamically. Notably, Huhns and Stephens [22] present a taxonomy of how agents coordinate behaviors and activities. In this framework, all elements (including cooperation and competition) emerge as forms of coordination. Coordination is seen as a system property, where agents operate within a shared environment, minimizing resource contention, avoiding livelock and deadlock, and maintaining safety conditions. In this taxonomy, cooperation is conceptualized as coordination among non-antagonistic agents, while negotiation represents coordination among competitive or self-interested agents. Then, successful cooperation typically would require agents to maintain models of one another and develop a model of future interactions, presupposing a degree of sociability.
@@ 3.3 Concept to Code Decay
As the system transitions from high-level concepts to operational structures, how can we ensure conceptual fidelity is maintained? Where are the translation gaps? What scaffolds (e.g., documentation, constraint encoding, behavior tracking) are in place to prevent conceptual fluctuation? Distortions may occur when theoretical models are translated into computational implementations, leading to unintended misalignments between the design intent and the real-world agents’ behavior. As a system transitions from high-level conceptual design to implementation and deployment, critical nuances may be altered, omitted, or misinterpreted. This often leads to a mismatch between initial goals (where cooperation may have been clearly defined) and deployed systems (where cooperation degrades into partial or only coordination). As a result, Concept-to-Code Decay is a key factor within the broader Misalignment Mosaic.
@@
Knowledge exists at multiple levels of abstraction [43], and abstraction is vital in shaping both high-level conceptualization and specific design decisions in MAS. Hence, the level of abstraction decreases as the system transitions from conceptual design to concrete implementation, while its representativeness of the real-world application increases, often improving its practical feasibility but sometimes at the cost of oversimplifying key functional and ethical considerations. This transition reminds of the Rabbit-Duck illusion: while a conceptual MAS model may capture intricate agent interactions and nuanced goals, its implementation may distort these complexities, either oversimplifying cooperative or coordinative mechanisms or introducing unintended emergent properties. This risks creating a gap between the system’s observable behaviors and its intended behavioral goals, which constitutes a form of misalignment. As a result, concept-to-code decay can lead to misalignment, where the system’s observable behaviors deviate from its intended design goals.
@@
Bădică and colleagues [44] inspire us to recognize that as abstraction decreases during the transition from conceptualization to deployment, key aspects of the original MAS motivation may be lost (concept-to-code decay). Addressing these abstraction gaps early in the design process is critical to ensuring that MAS implementations remain aligned with their intended objectives, minimizing both functional and ethical risks. Therefore, we must define terms and apply concepts carefully to align as much as possible within different stages, not only when informing a human about a system but also when designing, developing, testing, and assessing it. Misalignment can potentially lead to inefficiencies or even ethical concerns (especially when users are unaware of a system’s underlying decision-making paradigms).
@@
Figure 2: Illustration: A MAS transitioning from abstract concepts to concrete/operational.
@@
In Figure 2, we summarize the development and implementation of a MAS to illustrate the transition from abstract concepts to a concrete, operational entity. As the system progresses through stages – Conceptual Model, Setup/Design, Prototype, Testbed/Evaluation, and Deployment/Evolution – abstraction decreases while implementation details become more rigidly defined. However, as development advances, critical nuances (e.g., cooperative mechanisms, agent adaptability, and emergent behaviors) may be overlooked, leading to potential misalignment between the system’s observable behavior and its original design goals. While this example treats coordination and cooperation as emergent properties discovered during evaluation, developers could have embedded those as an explicit architectural choice early in the design phase. This flexibility underscores the importance of deliberate architectural decisions in shaping MAS behavior, ensuring that system alignment is proactively designed rather than emerging unpredictably.
@@ 3.4 Morality-As-Cooperation Theory
Achieving the best social outcome often requires agents to commit to cooperative strategies, i.e., agents must choose the actions that will only lead to the best outcome if most of the group commits and cooperates. However, despite its benefits, cooperation often results in a cost at the individual level, while defectors exploit shared resources [45]. This dilemma between self-interest and collective maintenance arises in strategic settings, such as the Prisoner’s Dilemma, and in real-world scenarios, including resistance to public health measures, where personal incentives undermine group stability.
@@
From an evolutionary perspective, cooperation plays a central role in structuring group cohesion and maintaining social systems [46]. Tomasello and Vaish [1] frame morality as an extension of cooperation, where pro-social motivations and cooperative behaviors scaffold the emergence of moral norms. Within this framing, cooperation requires alignment of individual and collective interests, often suppressing purely self-interested actions. Curry et al. [42] extend this link through a cross-cultural lens, analyzing how 60 societies conceptualize cooperation. They investigate seven cooperative behaviors (coordination to mutual advantage among them) to test the Morality-as-Cooperation Theory. Their findings suggest that studied behaviors are strong candidates for universal moral rules, reinforcing the idea that morality is rooted in cooperative principles. Thus, the Morality-As-Cooperation Theory highlights the notion that cooperation is deeply tied to moral rules, and understanding this connection is crucial for designing MAS that operate within ethical constraints.
@@
Besides, it makes us conceive cooperation in terms of sacrificial cooperation where self-interest or local utility is actively relinquished for the benefit of others, assuming no guaranteed reciprocity, and introducing loss, risk, or asymmetry into agent behavior, challenging utility-based models and exposing the limits of how cooperation is typically implemented in MAS. Thus, a question emerges: Does cooperation involve personal cost or risk without assured gain (sacrificial), or is it based on mutual benefit and rational strategy (strategic)?
@@
The distinction is crucial. In human societies, sacrificial cooperation (where individuals incur personal costs to uphold collective welfare, even absent guaranteed reciprocity) is key in moral systems [1, 45]. This raises fundamental questions about how MAS systems interpret, implement, and project cooperation, especially when deployed in environments where human observers may expect morally grounded, sacrificial behaviors — even when agents are only executing coordinative, utility-maximizing functions.
@@
Still, in the MAS literature, cooperation is frequently framed as both a desirable behavior and a technical goal. Yet this apparent neutrality conceals a deeper conflation: cooperation is often treated as a moral good, without acknowledging the ethical assumptions embedded in the term. In many MAS architectures, agents are labeled cooperative when they optimize for shared outcomes, but these outcomes rarely undergo moral scrutiny. The system appears trustworthy not because it embodies moral reasoning, but because the label cooperative carries inherent normative weight. This blurring of cooperation and morality creates an interpretive shortcut: agents designed for coordination or strategic alignment may be perceived as morally aligned, regardless of their internal architectures or value systems. Such overreadings have real consequences, shaping trust, system evaluation, and normative governance in ways that extend beyond technical performance.
@@
Recent work in the Foundations of Cooperative AI emphasizes these challenges [47]. Conitzer and Oesterheld frame cooperation as a theoretical cornerstone for alignment, grounded in game-theoretic models. While we agree that conceptual clarity is essential, our work highlights a preceding issue: without diagnosing meaning-level misalignment, efforts to formalize cooperation risk overlooking critical zones that precede behavioral misalignment. This Mosaic’s component (Morality-As-Cooperation Theory) urges the community to recognize when technical alignment gets unintentionally morally charged.

### v2

@@ 3 The Misalignment Mosaic
In MAS design, misalignment can occur not only from agent behavior but also from gaps between layers of design, implementation, and interpretation. Assumptions made early in the design process (often informal, idealized, or morally charged) get embedded in the architecture, translated into behavior, and eventually perceived by stakeholders as intentional system properties. This layered fluctuation creates a critical question: where exactly did the misalignment originate? And who is accountable for its consequences?
@@
Motivated by three interconnected drivers of meaning-level misalignment in MAS, the Misalignment Mosaic (depicted in Figure 1, and detailed in the following subsections) provides a structured diagnostic framework to trace how misalignment emerges from language, framing, and epistemic assumptions, beyond technical mechanisms alone. Rather than resolve ambiguity through formal redefinition, the Misalignment Mosaic enables audits of meaning, illuminating where alignment falters at the conceptual level.
@@
This framework builds on the concept of alignment magnification, which raises a critical question: how can AI systems be designed to pursue goals that remain consistent with their designers’ intentions? Dung [dung2023current] warns that as AI systems become more capable, misalignment risks increase, making these systems not only more dangerous but potentially harder to align. Therefore, alignment should be considered foundational in developing MAS: “No agent exists in a vacuum, and the evaluation of ethical behavior is a complex social and temporal phenomenon” [arnold2017value].
@@
In fact, the Morality-as-Cooperation Theory serves as a reminder that treating cooperation as a black box introduces a critical source of misalignment in MAS. This conflation is not just a semantic detail, but a structural risk factor in how agent behavior is interpreted, evaluated, and morally framed. Addressing alignment must therefore proceed in parallel with refining our understanding of coordination and cooperation (and their variants). The Misalignment Mosaic integrates seemingly distinct issues to show their role in challenging alignment in MAS. When system designers fail to address ambiguities, agents may operate under misaligned assumptions, leading to goal failures, inefficiencies, or emergent behaviors that contradict human intent. Clarifying these terms is essential for designing MAS systems that operate reliably within both technical and societal constraints.
To support semantic audits and reflective analysis, we are developing a Mosaic Reflection Sheet, an exploratory tool aligned with the Mosaic’s components. The approach of Winikoff et al. [winikoff2025scoresheet], who introduced a Scoresheet focused on explainability, inspired us. Their tool, applicable to MAS and broader AI technologies, incorporates stakeholder requirements and complements the IEEE P7001 Scoresheet [winfield2021ieee], which focuses on transparency. The Mosaic Reflection Sheet may include one reflection sheet for each of the four Misalignment Mosaic components, all designed to support interpretive clarity and trace epistemic commitments across MAS development stages.
Finally, our efforts contribute to recent Cooperative AI priorities by enabling audits of semantic assumptions and conceptual fluctuations, which are increasingly necessary as MAS operate in norm-sensitive and socially embedded environments. We detail the Mosaic’s four components next.
@@ 3.1 Terminological Inconsistency
The coordination/cooperation ambiguity leads to terminological inconsistencies, where terms are often defined interchangeably or imprecisely (see section 2). This fragmentation hinders theoretical and empirical progress, making it difficult to compare findings. Having cooperation still largely treated as a black box is no surprise, given that design patterns for MAS are not very popular.
A review of the design patterns literature [juziuk2014design] reveals a notable absence of standardized templates, significantly impacting MAS practical application. This lack of uniformity results in inconsistent implementations, misinterpretations, and reduced transferability across systems. Furthermore, according to Juziuk and colleagues [juziuk2014design], the relationships between patterns are poorly articulated, leading to a fragmented understanding of the pattern space: experts struggle to navigate and apply patterns effectively, limiting their practical utility. While MAS design patterns have been applied across various domains, existing classifications are largely confined to isolated catalogs, which prevents broader integration and adaptation [juziuk2014design]. A key example is the lack of standardized documentation and taxonomies for coordination and cooperation design patterns (along with variants). Without well-established standards, researchers apply similar terms to structurally distinct agent designs. While this issue also has interpretive consequences, we distinguish it from the Mosaic’s Interpretive Ambiguity by anchoring it in terminological and design-level fragmentation. As MAS applications extend to large-scale deployment across critical industries such as healthcare, transportation systems, and food and agriculture [keyIndustries], the demand for systems that are not only efficient but also transparent and ethically grounded intensifies.
@@ 3.2 Interpretive Ambiguity
Illustrated by the Rabbit-Duck illusion, this component highlights ambiguities and interpretational limits in MAS research. Just as perception shifts in the illusion, MAS frameworks can be framed differently depending on theoretical or empirical perspectives, leading to inconsistent definitions of coordination and cooperation. In complex or uncertain environments, incomplete knowledge and partial observability further blur the distinction between coordinating tasks and cooperating for collective benefit, thereby amplifying the reach of the Misalignment Mosaic. While the Rabbit-Duck illusion broadly represents perceptual ambiguity, we use it here to illustrate interpretive shift between cooperation and coordination (and related terms), a shift that has conceptual and moral consequences in MAS design.
Serious issues can arise when algorithmic systems’ goals are inferred solely from observations, potentially leading to misinterpretations of both the system’s behavior and its intended objectives. Thus, ambiguity in defining cooperation frequently emerges in practice: consider agricultural robots tasked with harvesting crops. At first glance, their interactions may appear cooperative since they collectively contribute to the same task. However, closer analysis may reveal only coordination, where robots follow predefined movement patterns to avoid overlap rather than actively adapting to each other’s presence or modifying strategies based on shared goals. Or, in another example, if various self-driving cars approach a four-way stop, should they coordinate by following predefined traffic rules or activate a cooperative mechanism that helps them dynamically yield? This ambiguity extends beyond system behavior to expert interpretation. Just as in the Rabbit-Duck illusion, it is possible that two researchers analyzing the same MAS testbed reach conflicting conclusions: one classifying it as a coordinative system, while the other sees cooperative behavior. This raises key epistemic and technical questions:
@@
As we try to address the questions above, we observe that multiple factors systematically influence whether MAS behavior is interpreted as coordinative or as indicative of cooperation. These factors span at least three levels of various considerations (note that [huhns1999multiagent] provides insights into those):
@@
3.2.1 If Coordination Exists, Is Cooperation Needed?
@@
Klassen et al. [klassen2023epistemic] demonstrate that partial knowledge can significantly impact agent beliefs and decision-making, sometimes leading to undesirable or even catastrophic outcomes. This raises a critical question for MAS: when agents face epistemic uncertainty, is coordination alone sufficient, or is cooperation necessary to mitigate these risks? In real-world deployments, agents rarely possess complete contextual knowledge due to limitations in data acquisition, processing, and environmental unpredictability.
While coordination alone may suffice in simpler, well-defined settings, real-world challenges require strategies that extend beyond simple coordination. Agents must navigate uncertainty, evolving conditions, and imperfect information, making cooperation a potentially necessary mechanism for adaptation and robustness. Given interpretive ambiguity, agents cannot rely solely on their knowledge and beliefs, as these limitations can adversely affect their decision-making capabilities (epistemic side effects [klassen2023epistemic]). This raises yet another question: if agents had perfect knowledge, would coordination alone be sufficient for optimal decision-making?
@@
While theoretically possible, agents will operate under knowledge constraints in practice. Given this reality, could a cooperative model serve as a secondary mechanism, allowing agents to compensate for knowledge gaps and make more reliable decisions? This aligns with Curry et al. [curry2019good], who demonstrate that cooperative frameworks, grounded in shared goals and moral principles, offer effective approaches for addressing uncertainty and achieving collective success. Incorporating cooperative mechanisms into MAS may thus provide a more effective means of addressing epistemic limitations, ensuring more resilient decision-making in uncertain environments.
@@
Focusing on distributed agent systems, Huhns and Stephens [huhns1999multiagent] treat agents collectively (as a society of agents) to address scalability challenges in environments with large agent populations. Their work presents a systematic approach to analyzing, describing, and designing multi-agent environments, taking into account both agent topology and content, which evolve dynamically. Notably, the authors [huhns1999multiagent] present a taxonomy of how agents coordinate behaviors and activities. In this framework, all elements (including cooperation and competition) emerge as forms of coordination. Coordination is seen as a system property, where agents operate within a shared environment, minimizing resource contention, avoiding livelock and deadlock, and maintaining safety conditions. In this taxonomy, cooperation is conceptualized as coordination among non-antagonistic agents, while negotiation represents coordination among competitive or self-interested agents. Then, successful cooperation typically would require agents to maintain models of one another and develop a model of future interactions, presupposing a degree of sociability.
A main takeaway from these considerations is that, to properly ask “If coordination exists, is cooperation needed?”, the MAS community needs a clearer and consistent distinction between coordination and cooperation.
@@ 3.3 Concept to Code Decay
As a MAS transitions from abstract concepts to operational structures, how can we ensure conceptual fidelity is maintained? Where are the translation gaps? What scaffolds (e.g., documentation, constraint encoding, behavior tracking) are in place to prevent conceptual fluctuation? Distortions may occur when theoretical models are translated into computational implementations, leading to unintended misalignments between the design intent and the real-world agents’ behavior. As a system transitions from high-level conceptual design to implementation and deployment, critical nuances may be altered, omitted, or misinterpreted.
This can lead to a mismatch between initial goals (where cooperation may have been clearly defined) and deployed systems (where cooperation degrades into partial or only coordination). As a result, Concept-to-Code Decay is a key factor within the broader Misalignment Mosaic.
@@
Knowledge exists at multiple levels of abstraction [abstractionTaxonomy], and abstraction is vital in shaping both high-level conceptualization and specific design decisions in MAS. Hence, the level of abstraction decreases as the system transitions from conceptual design to concrete implementation, while its representativeness of the real-world application increases, often improving its practical feasibility but sometimes at the cost of oversimplifying key functional and ethical considerations. This transition reminds us of the Rabbit-Duck illusion: while a conceptual MAS model may capture intricate agent interactions and nuanced goals, its implementation may distort these complexities, either oversimplifying cooperative or coordinative mechanisms or introducing unintended emergent properties. This risks creating a gap between the system’s observable behaviors and its intended behavioral goals, which constitutes a form of misalignment. As a result, concept-to-code decay can lead to misalignment, where the system’s observable behaviors deviate from its intended design goals.
@@
Bădică and colleagues [buadicua2018multi] inspire us to recognize that as abstraction decreases during the transition from conceptualization to deployment, key aspects of the original MAS motivation may be lost (concept-to-code decay). Addressing these abstraction gaps early in the design process is critical to ensuring that MAS implementations remain aligned with their intended objectives, minimizing both functional and ethical risks. Therefore, we must define terms and apply concepts carefully to align as much as possible within different stages, not only when informing a human about a system but also when designing, developing, testing, and assessing it. Misalignment can potentially lead to inefficiencies or even ethical concerns (especially when users are unaware of a system’s underlying decision-making paradigms).
@@
Figure 2: An illustration of a MAS transitioning from abstract concepts to concrete/operational.
@@
In Figure 2, we summarize the development and implementation of a MAS to illustrate the transition from abstract concepts to a concrete, operational entity. As the system progresses through stages – Conceptual Model, Setup/Design, Prototype, Testbed/Evaluation, and Deployment/Evolution – abstraction decreases while implementation details become more rigidly defined. However, as development advances, critical nuances (e.g., cooperative mechanisms, agent adaptability, and emergent behaviors) may be overlooked, leading to potential misalignment between the system’s observable behavior and its original design goals. While this example treats coordination and cooperation as emergent properties discovered during evaluation, developers could have embedded them as explicit architectural choices early in the design phase. This flexibility underscores the importance of deliberate architectural decisions in shaping MAS behavior, ensuring that system alignment is proactively designed rather than emerging unpredictably. On a side note, Norval and colleagues [norval2022disclosure] offer interesting insights into document engineering, while [ElFassi02102024] into assumptions using change propagation as context (they also reflect on assumption definition); finally, UX/UI design also provides insights into reflecting on design assumptions.
@@ 3.4 Morality-As-Cooperation Theory
Achieving the best social outcome often requires agents to commit to cooperative strategies, i.e., agents must choose the actions that will only lead to the best outcome if most of the group commits and cooperates. However, despite its benefits, cooperation often results in a cost at the individual level, while defectors exploit shared resources [wardil2014origin]. This dilemma between self-interest and collective maintenance arises in strategic settings, such as the Prisoner’s Dilemma, and in real-world scenarios, including resistance to public health measures, where personal incentives undermine group stability.
@@
From an evolutionary perspective, cooperation plays a central role in structuring group cohesion and maintaining social systems [tomasello2011human]. Tomasello and Vaish [tomasello2013origins] frame morality as an extension of cooperation, where pro-social motivations and cooperative behaviors scaffold the emergence of moral norms. Within this framing, cooperation requires alignment of individual and collective interests, often suppressing purely self-interested actions. Curry et al. [curry2019good] extend this link through a cross-cultural lens, analyzing how 60 societies conceptualize cooperation. They investigate seven cooperative behaviors (coordination to mutual advantage among them) to test the Morality-as-Cooperation Theory. Their findings suggest that studied behaviors are strong candidates for universal moral rules, reinforcing the idea that morality is rooted in cooperative principles. Thus, the Morality-As-Cooperation Theory highlights the notion that cooperation is deeply tied to moral rules, and understanding this connection is crucial for designing MAS that operate within ethical constraints.
@@
Besides, it calls attention to cooperation defined in terms of sacrificial cooperation where self-interest or local utility is actively relinquished for the benefit of others, assuming no guaranteed reciprocity, and introducing loss, risk, or asymmetry into agent behavior, challenging utility-based models and exposing the limits of how cooperation is typically implemented in MAS. Thus, a question emerges: does cooperation involve individual costs or risks without assured gain (sacrificial), or is it based on mutual benefit and rational strategy (strategic)?
@@
The distinction is crucial. In human societies, sacrificial cooperation (where individuals incur personal costs to uphold collective welfare, even absent guaranteed reciprocity) is central in moral systems [tomasello2013origins, wardil2014origin]. This raises fundamental questions about how MAS systems interpret, implement, and project cooperation, especially when deployed in environments where human observers may expect morally grounded, sacrificial behaviors — even when agents are only executing coordinative, utility-maximizing functions.
@@
Still, in the MAS literature, cooperation is frequently framed as both a desirable behavior and a technical goal. Yet this apparent neutrality hides a deeper conflation: cooperation is often treated as a moral good, without acknowledging the ethical assumptions embedded in the term. In many MAS architectures, agents are labeled cooperative when they optimize for shared outcomes, but these outcomes rarely undergo moral scrutiny. The system appears trustworthy not because it embodies moral reasoning, but because the label cooperative carries inherent normative weight.
This blurring of cooperation and morality creates an interpretive shortcut: agents designed for coordination or strategic alignment may be perceived as morally aligned, regardless of their internal architectures or value systems. Such overreadings have real consequences, shaping trust, system evaluation, and normative governance in ways that extend beyond technical performance.
@@
Recent work in the Foundations of Cooperative AI highlights these challenges [conitzer2023foundations]. Conitzer and Oesterheld frame cooperation as a theoretical cornerstone for alignment, grounded in game-theoretic models. While we agree that conceptual clarity is essential, our work highlights a preceding issue: without diagnosing meaning-level misalignment, efforts to formalize cooperation risk overlooking critical zones that precede behavioral misalignment. This Mosaic’s component (Morality-As-Cooperation Theory) urges the community to recognize when technical alignment gets unintentionally morally charged.